OpenStudy (anonymous):

how many liters (L) of pure water should be mixed with a 5-L solution of 60% acid to produce a mixture that is 50% water?

OpenStudy (anonymous):

can I help

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I am here

OpenStudy (anonymous):

how much present an% do you think

OpenStudy (anonymous):

@Jones2015Calvin

OpenStudy (anonymous):

It's asking how many liters

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

What are you asking me?

OpenStudy (anonymous):

common

OpenStudy (anonymous):

how much present you know the answer

OpenStudy (anonymous):

means what do you think the answere is

OpenStudy (anonymous):

Would it be set up as 0x+.6(5)= .5(5+x)

OpenStudy (anonymous):

noooooooooo

OpenStudy (anonymous):

Then I don't know the answer

OpenStudy (anonymous):

I under stand that's why I am here for

OpenStudy (anonymous):

So how should it be set up ?

OpenStudy (wolf1728):

5 L of 60% acid =3 L of acid 2 L of water How much water to add to make it 50% water? concentration = acid volume / total volume .6 = 3L / 5L We need .50 Concentration so .5 = 3L / Total Volume Total Volume = 3L / .5 Total Volume is 6 L so we must add 1 liter of water

OpenStudy (anonymous):

10

OpenStudy (wolf1728):

Why do you say 10?

OpenStudy (anonymous):

Difference between.5 and.6

OpenStudy (wolf1728):

I'm pretty sure it is 1 liter. Did you look at my answer?

OpenStudy (anonymous):

I see it now

OpenStudy (wolf1728):

So, I'm right?

OpenStudy (anonymous):

Yes now it's asking how many Liters of pure water should be mixed with a 6-L solution of 90% acid to produce a mixture that 40% water

OpenStudy (wolf1728):

6L of 90% acid How much water to add to make it 40% water? 6L .90 acid = 5.4 L acid .6 L water We need 40% water so concentration = acid volume / total volume .4 = 5.4 L / total volume total volume = 5.4L / .4 total volume = 13.5 L So, we have to add 13.5 - 6L = 7.5 L water

OpenStudy (anonymous):

Where the 5.4 come from

OpenStudy (wolf1728):

The original solution is 6 L of 90% acid So, original solution = 90% acid 10% water acid volume = .9 * 6 = 5.4 Liters water volume = .1 * 6 = .6 liters

OpenStudy (anonymous):

What's next

OpenStudy (perl):

.

OpenStudy (anonymous):

Can you give me a better understanding

OpenStudy (qwertty123):

From the view of acid it is since 100% water has 0% acid 0*x+.6*8=0.5(8+x) 4.8=4+.5x .8=.5x .8/.5=x x=1.6 L view of water I changed from percent acid to percent non acid 60% acid means 40 % non acid 50% acid means 50 % non acid

OpenStudy (qwertty123):

@Jones2015Calvin

OpenStudy (anonymous):

Would that work for how many Liters of pure water should be mixed with a 6-L solution of 90% acid to produce a mixture that 40% water

OpenStudy (qwertty123):

Pure water is 100% non acid as a decimal 100%=1 1*x+0.4*8=0.5(8+x) 1*x+3.2=4+0.5x 1x-0.5x=4-3.2 0.5x=0.8 x=1.6 L Same answer!!!! check 1*1.6+0.4*8=0.5(8+1.6) 1.6+3.2=0.5(9.6) 4.8=4.8 ok

OpenStudy (qwertty123):

Oh your Equation: water + water = water 1*x + 0.40*8 = 0.50(x+8)

OpenStudy (qwertty123):

@Jones2015Calvin Is that another question?

OpenStudy (wolf1728):

I guess Jones2015Calvin isn't here anymore. I'm surprised h still remained. I left here about an hour ago.

OpenStudy (qwertty123):

oh lol

OpenStudy (anonymous):

Yes I'm here